# Coherent Multi-Agent Trajectory Forecasting in Team Sports with CausalTraj

### v1

@@ Introduction
Artificial intelligence is increasingly applied in sports analytics, enabling automatic understanding of tactical patterns, player roles, and decision-making dynamics (tuyls2021game; wang2024tacticai; wu2022sports). A key foundation for such analysis is trajectory prediction – forecasting the future motion of players and the ball from historical observations. Accurate trajectory forecasting not only aids tactical interpretation but also enables realistic simulation and generation of gameplay sequences. Beyond sports, this problem also appears in autonomous driving (8793868) and crowd navigation (Lisotto_2019_ICCV), where interaction and coordination among multiple agents are central.
@@
Trajectory forecasting in these domains is inherently challenging: the future is stochastic and multimodal, and each agent’s motion depends on the collective configuration of all others. We observe that most recent state-of-the-art works on sports datasets (mao2023leapfrog; lee2024mart; fu2025moflowonestepflowmatching) design their models and objectives to be optimal for per-agent metrics such as minADE or minFDE. These metrics evaluate each agent independently on its best-of-k predicted trajectory, which encourages models to optimize marginal accuracy, often by increasing sample diversity, rather than learning how trajectories should fit together as a joint multi-agent future. As a result, models may achieve good per-agent scores yet still underperform on joint prediction. They may also generate trajectories that look reasonable in isolation but fail to form coherent  [^1] We use ”coherent” to refer to a property of multi-agent trajectories in which agents’ motions are physically plausible and mutually compatible, e.g. ball motion aligning with player motion and players positioning themselves in coordinated formations. multi-agent evolutions. The importance of assessing joint modelling capability via joint accuracy metrics has formerly been highlighted in other domains involving group dynamics (weng2023jointmetricsmatterbetter). In team sports, the state of play is defined by joint behaviors. As we cannot assume oracle knowledge of groundtruth to select compatible marginal predictions at test time, reliable joint predictions are fundamental.
@@
We view the capability to learn the true joint distribution as the central goal to pursue. When the true joint distribution is captured, scenario-level coherence and strong per-agent accuracy should emerge naturally as byproducts rather than competing objectives. Inspired by recent successes of causal (autoregressive) architectures in language (LLMs) and 3D environment generation  (radford2019language; parkerholder2024genie2), we revisit the temporally causal framework for trajectory prediction. Modeling causality in time allows spatial and inter-agent dynamics to evolve step-by-step, rather than being compressed into a fixed latent representation. Combined with likelihood-based objectives, it captures multimodal transition uncertainty while providing high capacity for joint modeling.
@@
Causal (autoregressive) formulations have shown success in earlier works (IvanovicP19; zhan2019generating). However, many recent state-of-the-art methods achieving top per-agent metrics on sports datasets adopt designs that predict the full future trajectory horizon in parallel (mao2023leapfrog; lee2024mart; fu2025moflowonestepflowmatching). This formulation, often paired with multi-branch or multi-sample output heads, enables explicit control over prediction diversity and simplifies optimization for marginal (per-agent) accuracy. However, for joint trajectory predictions, predicting all agents and timesteps simultaneously imposes strong conditional-independence assumptions on the shared latent representation. As a result, modeling long-horizon interdependent agent dynamics would probably require a huge and expressive latent state.
@@
We revisit the causal formulation, motivated by its proven success in capturing high-order dependencies across sequential domains such as language and 3D environment generation (radford2019language; parkerholder2024genie2). By modeling interactions step-by-step instead of compressing them into a latent state, the causal structure reduces latent capacity demands. With lowered stochastic capacity requirement, we directly parameterize multimodal trajectory likelihoods in the output space, enabling exact optimization of probabilistic objectives without approximate inference. Our architecture combines a transformer-based inter-agent relation encoder with spatial adaptations and a Mamba2-based temporal encoder under a unified causal framework. By integrating these complementary components, CausalTraj produces coherent multi-agent forecasts that excel on joint metrics.

### v2

@@ Introduction
Artificial intelligence is increasingly applied in sports analytics, enabling automatic understanding of tactical patterns, player roles, and decision-making dynamics (tuyls2021game; wang2024tacticai; wu2022sports). A key foundation for such analysis is trajectory prediction – forecasting the future motion of players and the ball from historical observations. Accurate trajectory forecasting not only aids tactical interpretation but also enables realistic simulation and generation of gameplay sequences. Beyond sports, this problem also appears in autonomous driving (8793868) and crowd navigation (Lisotto_2019_ICCV), where interaction and coordination among multiple agents are central. Trajectory forecasting in these domains is inherently challenging: the future is stochastic and multimodal, and each agent’s motion depends on the collective configuration of all others.
@@
Most trajectory forecasting models are evaluated solely using per-agent metrics (minADE and minFDE), which score each agent independently based on its best-of-k predicted trajectories. Many state-of-the-art works on sports datasets (mao2023leapfrog; lee2024mart; fu2025moflowonestepflowmatching) design their models and training objectives around these metrics. For instance, their losses supervise multi-hypothesis trajectory selection independently for each agent, without modeling which predicted trajectories across agents should fit together to form a joint multi-agent future. As a result, models may achieve good per-agent scores yet still underperform on joint prediction. They may also generate trajectories that look reasonable individually but fail to form coherent  [^1] We use ”coherent” to refer to a property of multi-agent trajectories in which agents’ relative motions are physically plausible and mutually compatible, e.g. ball motion aligning with player motion and players positioning themselves in coordinated formations. multi-agent evolutions. The importance of assessing joint predictions in trajectory forecasting has formerly been highlighted in other domains involving group dynamics (weng2023jointmetricsmatterbetter). In team sports, the state of play is defined by joint behaviors. As we cannot assume oracle knowledge of groundtruth to select compatible marginal predictions at test time, reliable joint predictions are fundamental.
@@
We view the capability to learn the true joint distribution as the central goal to pursue. When the true joint distribution is captured, scenario-level coherence and strong per-agent accuracy should emerge naturally as byproducts rather than competing objectives. Inspired by recent successes of causal (autoregressive) architectures in language (LLMs) and 3D environment generation  (radford2019language; parkerholder2024genie2), we revisit the temporally causal framework for trajectory prediction. Modeling causality in time allows spatial and inter-agent dynamics to evolve step-by-step, rather than being compressed into a fixed global latent representation. Combined with likelihood-based objectives, it captures multimodal transition uncertainty while providing high capacity for joint modeling.
@@
Causal (autoregressive) formulations have shown success in earlier works (IvanovicP19; zhan2019generating). However, many recent state-of-the-art methods achieving top per-agent metrics on sports datasets adopt designs that predict the full future trajectory horizon in parallel (mao2023leapfrog; lee2024mart; fu2025moflowonestepflowmatching), often via an intermediate global latent representation. This formulation, paired with multi-sample output heads, enables control over prediction diversity and simplifies optimization for marginal (per-agent) accuracy. However, for joint trajectory predictions, predicting all agents and timesteps simultaneously require the outputs across timesteps to be conditionally independent given the global latent representation. As a result, modeling interdependent agent dynamics over a long horizon would probably require a huge and expressive latent state.
@@
We revisit the causal formulation, motivated by its proven success in capturing high-order dependencies across sequential domains such as language and 3D environment generation (radford2019language; parkerholder2024genie2). By modeling interactions step-by-step instead of compressing them into a latent state, the causal structure reduces the requirement for latent capacity. We directly parameterize multimodal trajectory likelihoods in the output space, enabling exact optimization of probabilistic objectives without approximate inference. Our architecture combines a transformer-based inter-agent relation encoder with spatial adaptations and a Mamba2-based temporal encoder under a unified causal framework. By integrating these complementary components, CausalTraj produces coherent multi-agent forecasts that excel on joint metrics.